OpenStudy (anonymous):

How would you integrate - \[\int\limits_{?}^{?} (\sin x / \sin 4x) dx\] ?

OpenStudy (anonymous):

It's an indefinite integral.

OpenStudy (anonymous):

Maybe you can use \[\sin(4x) = 8\cos(x)^3\sin(x) - 4\cos(x)\sin(x) \]

OpenStudy (anonymous):

I did that. (And a lot more actually.) But the problem is that if I substitute cos x = y (say), I am left with an extra term and the integral becomes very very complicated!

OpenStudy (anonymous):

Tried integration by parts where \[ u=\frac{1}{\sin(4x)} \] ?

OpenStudy (anonymous):

That gives you the integral of cot(x)^2 which is easy because \[ \frac{d}{dx} \cot(x) = -1-\cot(x)^2 \]

OpenStudy (anonymous):

Do you see it or should I write it out for you?

OpenStudy (anonymous):

Thanks I got it!